Using historical records, the personnel manager of a plant has determined the probability of XX, the number of employees absent per day. It is

X 0 1 2 3 4 5 6 7
P(X) 0.0046 0.0248 0.3098 0.3399 0.219 0.0798 0.019 0.0031

Find the following probabilities.

A. P(2≤X≤5)P(2≤X≤5)

Probability =

B. P(X>5)P(X>5)

Probability =

C.P(X<4)P(X<4)

Probability =
Answer #1

a)

P(2 <= X <= 5) = P(X = 2) + P(X = 3) + P(X = 4) + P(X = 5)

= 0.3098 + 0.3399 + 0.2190 + 0.0798

= 0.9485

b)

P(X > 5) = P(X >= 6)

= P(X = 6) + P(X = 7)

= 0.019 + 0.0031

= 0.0221

c)

P(X < 4) = P(X <= 3)

= P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3)

= 0.0046 + 0.0248 + 0.3098 + 0.3399

= 0.6791